OpenStudy (anonymous):

im trying to find the limit of 1.001^x as x approaches negative infinity. i know the answer is 0 because 1.001 raised to an decreasing negative number gives a value less than zero. but i wanted to know how would you show the work to prove that equal zero rather than just writing that the answer is zero

OpenStudy (anonymous):

By "value less than zero" I think you mean either "value between 1 and 0" or something along the lines of a "value getting arbitrarily small and approaching 0." If you're looking for a somewhat rigorous approach to establishing that the limit is indeed 0 as \(x\to-\infty\), you can focus on the sequence \(a_n=\dfrac{1}{1.001^n}\) and show that \(a_n\) is both decreasing and bounded below. You can establish these facts via induction. To prove the claim that \(a_n\) is decreasing, you want to show that \(a_n>a_{n+1}\) for all \(n\ge1\). The basis case, that \(a_1<a_2\), is clear, since \(\dfrac{1}{1.001}>\dfrac{1}{1.001^2}\) because \(1.001<1.001^2\). Assume this relation holds for \(n=k\), i.e. that \(a_k>a_{k+1}\), and use this to show that the relation holds for \(n=k+1\), i.e. that \(a_{k+1}>a_{k+2}\). \[a_{k+2}=\frac{1}{1.001^{k+2}}=\frac{1}{1.001}\times\frac{1}{1.001^{k+1}}<\frac{1}{1.001}\times\frac{1}{1.001^k}=\frac{1}{1.001^{k+1}}=a_{k+1}\] So, \(a_k>a_{k+1}\) implies \(a_{k+1}>a_{k+2}\), and so \(a_n\) is a decreasing sequence. Boundedness is a direct result of the fact that \(a^b>0\) for any real \(a>0\), \(b\neq0\), so \(a_n=\dfrac{1}{1.001^n}>0\) for all \(n\). Taken together, the fact that the sequence is bounded below by 0 AND is decreasing, you can conclude that the limit of the sequence is 0, which proves the limit of your function is also 0. (The sequence \(a_n\) is essentially the discrete form of your function.)